# ISAC Network Planning: Sensing Coverage Analysis and 3-D BS Deployment Optimization

### v1

@@ V-B Sensing Performance of 1D Area
To contrast to BS deployments optimized for area-sensing performance to those tailored for communication coverage, we first consider an elementary one-dimensional user region along the line $y=0$, $x\in[0,1000]$. Users (green rectangle) are placed uniformly along this line, and a fixed number of BSs (red triangles) are placed according to two distinct strategies. Targets (blue dots) are likewise interspersed along the same line to evaluate sensing performance. See Figs. 5, 7, 9, and 10 for illustration. In Fig. 5(a)), all BSs lie as close as possible to the communication user line, i.e., above $y=0$ with their $x$–coordinates evenly spaced over $[0,1000]$. This arrangement minimizes the path-loss for each BS-user link, yielding maximal end-to-end SNR for purely communication-centric operation. By contrast, in Fig. 5(b)), to balance the associated sensing signal power and geometry gain requirements, two BSs remain near the target line (at $y\approx 0$) to preserve low-attenuation links, while the remaining BSs are shifted in $y$ to form near-perpendicular baselines relative to the target line. This bi-static geometry, yielding approximately $90^{\circ}$ incidence angles at most targets enhances geometric gain, thereby improving area-wide sensing accuracy without unduly sacrificing communication performance.
@@
Fig. 6 shows the minimal CRLB as a function of the one-dimensional region length $L$ and various BS heights. In each curve, the region length is increased from 250m to 1000m, and the optimal BS height is selected to minimize the averaged CRLB over all target positions in the interval $[0,L]$. Moreover, when $L$ is multiplied by an integer factor, the optimal BS height $h^{*}$ also increases by roughly the same factor. This one-to-one proportionality between region size and BS height is a direct consequence of the geometric invariance properties. Explicitly, scaling the coverage interval by $\kappa$ only preserves the angular geometry if the BS elevation is scaled by $\kappa$ as well, as analyzed in Proposition 4. The dominant effect driving the optimal BS height is the geometric gain in target sensing. Increasing the number of BSs in proportion to the region length maintains consistent incidence angles and range spreads, thereby balancing the multiple impact factors in the CRLB across the entire coverage area. These numerical findings validate our theoretical derivations and provide concrete guidelines for selecting the BS heights in large-scale ISAC deployments.
@@ V-C Sensing Performance of 2D Area
As illustrated in Fig. 7, we assess the two-dimensional coverage performance of communication and sensing services, where the communication users are uniformly distributed on the ground plane (green) and targets over a parallel plane at an altitude of 400m (blue). Figs. 7(a) and 7(d), 7(b) and 7(e), as well as 7(c) and 7(f) depict three deployment strategies maximizing communication rate, minimizing the CRLB under a 5 bps/Hz communication constraint, and minimizing the CRLB without any communication requirement, respectively. In the communication-only scenario (c.f. Fig. 7(a) and Fig. 7(d)), BSs are arranged symmetrically across the area, yielding a symmetric CRLB map with values ranging from $10^{-1.2}$ to $110^{0}=$ 1. When enforcing the 5 bps/Hz rate (c.f. 7(b) and 7(e)), the BSs shift slightly toward the sensing plane and adjust their planar geometry to balance both objectives. Without any communication constraint (c.f. 7(c) and 7(f)), the optimal sensing configuration moves BSs closer to the target layer and reconfigures their layout to minimize the CRLB, where even a modest height adjustment of a single BS combined with a small horizontal displacement yields significant sensing performance gains. Compared to the communication-optimal deployment, the overall positioning CRLB map is improved by a factor of $4$. In Fig. 7(c), it is observed that simply minimizing the distance between BSs and the sensing region tends to place them as close as possible to the targets. However, while a shorter distance increases the scaling factor, it also reduces the angular diversity among the BSs. This causes the row vectors of the Jacobian matrix to become nearly linearly dependent, drastically increasing the condition number of the Fisher information matrix. To balance the angular spread and measurement accuracy, BSs should be deployed at an appropriate distance from the sensing region, which is similar to the results in Fig. 5(b).
@@
(c) Sensing-optimal CRLB map.
@@
(f) Sensing-optimal CRLB map.
@@
Fig. 9(a) considers a conventional cellular-style layout in which each BS is placed on a regular grid (horizontal coordinates fixed) at a uniform height of 10m. Despite proximity to some BSs, severe sensing-coverage voids remain, particularly in regions where the geometric intersection angles of the sensing links are acute, a consequence of BSs being deployed relatively close to the target plane. To isolate the impact of antenna elevation, we then retain the same horizontal BS locations but optimize only their heights. As shown in Fig. 9(b), this vertical repositioning compresses the CRLB values from the original span of $[10^{-3},10^{2}]$ down to $[10^{-5},10^{0}]$, reflecting a dramatic gain in localization precision. Finally, by applying a CRLB threshold of $10^{-1}$, Fig. 9(c) demonstrates that sensing coverage jumps from roughly 10% to 60%, illustrating that even modest BS-height adjustments can yield substantial improvements in network-level ISAC performance.
In Fig. 10, we compare the BS layouts optimized for various spatial probability density functions (PDF) of sensing targets and communication users. Let the service area be denoted by $\mathcal{A}$, and assume that the user and target locations follow the known density functions $P_{u}(\mathbf{u})$ and $P_{t}(\mathbf{t})$, respectively, which can be estimated a priori from historical access logs, traffic heatmaps, or mobility statistics. Under these distributions, the optimal BSs naturally cluster around high-density regions to enhance sensing performance. However, to avoid poor geometric intersection angles that degrade localization accuracy, the deployment must also include at least one elevated pivot BS positioned near the overall centroid of $\mathcal{A}$. This hybrid strategy both preserves fairness across the service area and significantly improves geometric gain for multi-static sensing. Consequently, areas having higher user or target concentration benefit from enhanced sensing coverage, while the pivot BS ensures robust performance in sparsely populated regions.
@@
As illustrated in Fig. 11, optimizing the BS deployment purely for maximum communication throughput yields a peak rate of 5.75 bps/Hz, but results in uniformly low sensing coverage $\Pr(\mathrm{CRLB}\leq\Gamma)$ across all CRLB thresholds. By imposing a modest throughput constraint, reducing the peak rate to 5 bps/Hz (a 23% decrease), we achieve roughly a 20% uplift in sensing coverage across the board. Finally, if the optimization objective is switched to minimizing the CRLB with no communication-rate requirement, the sensing coverage at $\Gamma=0.1$ increases by nearly threefold. These findings vividly illustrate the inherent trade-off between communication throughput and sensing performance when designing BS deployments. Fig. 11 illustrates how base-station placement governs the trade-off between communication and sensing performance. As the minimum communication rate requirement is tightened, sensing coverage deteriorates, underscoring the influence of BS geometry on this dual-function trade-off. Remarkably, by perturbing the positions of just a few BSs, at the cost of under a 10% reduction in peak data rate, sensing performance can be boosted substantially. Compared to the baseline layout optimized solely for maximum throughput, these minor relocations yield a network configuration that simultaneously satisfies both communication and sensing criteria. Moreover, Fig. 11 reveals that this trade-off exacerbates at higher target altitudes. Explicitly, as the mean separation between the sensing target region and the communication user region increases, maintaining both objectives simultaneously becomes increasingly difficult.
@@ VI Conclusions
A novel deployment framework was proposed for ISAC networks that addresses the fundamental tradeoff between S&C. Our design ensures both localization coverage and communication performance, guaranteeing high-precision localization and high-throughput data transmission across the entire service region. We analyzed the complexity of cooperative sensing among multiple BSs from a localization-performance coverage perspective, ensuring that the CRLB requirements are met uniformly across the region. By identifying key invariance properties of the optimal deployment, such as shift, rotation, and scaling invariance, we developed a low-complexity algorithmic framework for ISAC network planning. We formulated the joint sensing-communication optimization as a structured problem amenable to the MM principle, resulting in an MM-based algorithm that reliably converges to high-quality solutions at a low computational cost per iteration. Simulation results demonstrated that the proposed framework significantly enhances sensing coverage, while maintaining communication throughput with minimal performance degradation. This work provides insights for designing practical large-scale ISAC networks.

### v2

@@ V-B Sensing Performance of 1D Area
To contrast BS deployments optimized for area‐sensing performance to those tailored for communication coverage, we first consider an elementary one‐dimensional user region along the line $y=0$, $x\in[0,1000]$. Users (green rectangle) are placed uniformly along this line, and a fixed number of BSs (red triangles) are placed according to two distinct strategies. Targets (blue dots) are likewise interspersed along the same line to evaluate sensing performance. See Figs. 5, 7, 9, and 10 for illustration. In Fig. 5(a)), all BSs lie as close as possible to the communication user line, i.e., above $y=0$ with their $x$-coordinates evenly spaced over $[0,1000]$. This arrangement minimizes the path-loss for each BS-user link, yielding maximal end‐to‐end SNR for purely communication-centric operation. By contrast, in Fig. 5(b)), to balance the associated sensing signal power and geometry gain requirements, two BSs remain near the target line (at $y\approx 0$) to preserve low‐attenuation links, while the remaining BSs are shifted in $y$ to form near‐perpendicular baselines relative to the target line. This bi‐static geometry, yielding approximately $90^{\circ}$ incidence angles at most targets enhances geometric gain, thereby improving area‐wide sensing accuracy without unduly sacrificing communication performance. As shown in Fig. 5(c), under the communication constraint the optimal deployment places several BSs near the user region to boost link SNR, while positioning one BS at higher altitude to enhance sensing geometry over the target area, i.e., to increase angular diversity.
@@
Fig. 6 shows the minimal CRLB as a function of the one‐dimensional region length $L$ and various BS heights. In each curve, the region length is increased from 250m to 1000m, and the optimal BS height is selected to minimize the averaged CRLB over all target positions in the interval $[0,L]$. Moreover, when $L$ is multiplied by an integer factor, the optimal BS height $h^{*}$ also increases by roughly the same factor. This one‐to‐one proportionality between region size and BS height is a direct consequence of the geometric invariance properties. Explicitly, scaling the coverage interval by $\kappa$ only preserves the angular geometry if the BS elevation is scaled by $\kappa$ as well, as analyzed in Proposition 4. The dominant effect driving the optimal BS height is the geometric gain in target sensing. Increasing the number of BSs in proportion to the region length maintains consistent incidence angles and range spreads, thereby balancing the multiple impact factors in the CRLB across the entire coverage area. These numerical findings validate our theoretical derivations and provide concrete guidelines for selecting the BS heights in large‐scale ISAC deployments.
@@ V-C Sensing Performance of 2D Area
As illustrated in Fig. 7, we assess the two‐dimensional coverage performance of communication and sensing services, where the communication users are uniformly distributed on the ground plane (green) and targets over a parallel plane at an altitude of 400m (blue). Figs. 7(a) and 7(d), 7(b) and 7(e), as well as 7(c) and 7(f) depict three deployment strategies maximizing communication rate, minimizing the CRLB under a 5 bps/Hz communication constraint, and minimizing the CRLB without any communication requirement, respectively. In the communication‐only scenario (c.f. Fig. 7(a) and Fig. 7(d)), BSs are arranged symmetrically across the area, yielding a symmetric CRLB map with values ranging from $10^{-1.2}$ to $10^{0}=1$. When enforcing the 5 bps/Hz rate (c.f. 7(b) and 7(e)), the BSs shift slightly toward the sensing plane and adjust their planar geometry to balance both objectives. Without any communication constraint (c.f. 7(c) and 7(f)), the optimal sensing configuration moves BSs closer to the target layer and reconfigures their layout to minimize the CRLB, where even a modest height adjustment of a single BS combined with a small horizontal displacement yields significant sensing performance gains. Compared to the communication‐optimal deployment, the overall positioning CRLB map is improved by a factor of $4$. In Fig. 7(c), it is observed that simply minimizing the distance between BSs and the sensing region tends to place them as close as possible to the targets. However, while a shorter distance increases the scaling factor, it also reduces the angular diversity among the BSs. This causes the row vectors of the Jacobian matrix to become nearly linearly dependent, drastically increasing the condition number of the Fisher information matrix. To balance the angular spread and measurement accuracy, BSs should be deployed at an appropriate distance from the sensing region, which is similar to the results in Fig. 5(b).
@@
(c) Sensing‐optimal CRLB map.
@@
(f) Sensing‐optimal CRLB map.
@@
Fig. 9(a) considers a conventional cellular‐style layout in which each BS is placed on a regular grid (horizontal coordinates fixed) at a uniform height of 10m. Despite proximity to some BSs, severe sensing‐coverage voids remain, particularly in regions where the geometric intersection angles of the sensing links are acute, a consequence of BSs being deployed relatively close to the target plane. To isolate the impact of antenna elevation, we then retain the same horizontal BS locations but optimize only their heights. As shown in Fig. 9(b), this vertical repositioning compresses the CRLB values from the original span of $10^{-3},10^{2}]$ down to $10^{-5},10^{0}]$, reflecting a dramatic gain in localization precision. Finally, by applying a CRLB threshold of $10^{-1}$, Fig. 9(c) demonstrates that sensing coverage jumps from roughly 10% to 60%, illustrating that even modest BS‐height adjustments can yield substantial improvements in network‐level ISAC performance.
In Fig. 10, we compare the BS layouts optimized for various spatial probability density functions (PDF) of sensing targets and communication users. Let the service area be denoted by $\mathcal{A}$, and assume that the user and target locations follow the known density functions $P_{u}(\mathbf{u})$ and $P_{t}(\mathbf{t})$, respectively, which can be estimated a priori from historical access logs, traffic heatmaps, or mobility statistics. Under these distributions, the optimal BSs naturally cluster around high-density regions to enhance sensing performance. However, to avoid poor geometric intersection angles that degrade localization accuracy, the deployment must also include at least one elevated pivot BS positioned near the overall centroid of $\mathcal{A}$. This hybrid strategy both preserves fairness across the service area and significantly improves geometric gain for multi‐static sensing. Consequently, areas having higher user or target concentration benefit from enhanced sensing coverage, while the pivot BS ensures robust performance in sparsely populated regions.
@@
As illustrated in Fig. 11, optimizing the BS deployment purely for maximum communication throughput yields a peak rate of 5.75 bps/Hz, but results in uniformly low sensing coverage $\Pr(\mathrm{CRLB}\leq\Gamma)$ across all CRLB thresholds. By imposing a modest throughput constraint, reducing the peak rate to 5 bps/Hz (a 23% decrease), we achieve roughly a 20% uplift in sensing coverage across the board. Finally, if the optimization objective is switched to minimizing the CRLB with no communication‐rate requirement, the sensing coverage at $\Gamma=0.1$ increases by nearly threefold. These findings vividly illustrate the inherent trade‐off between communication throughput and sensing performance when designing BS deployments. Fig. 11 illustrates how base-station placement governs the trade-off between communication and sensing performance. As the minimum communication rate requirement is tightened, sensing coverage deteriorates, underscoring the influence of BS geometry on this dual-function trade-off. Remarkably, by perturbing the positions of just a few BSs, at the cost of under a 10% reduction in peak data rate, sensing performance can be boosted substantially. Compared to the baseline layout optimized solely for maximum throughput, these minor relocations yield a network configuration that simultaneously satisfies both communication and sensing criteria. Moreover, Fig. 11 reveals that this trade-off exacerbates at higher target altitudes. Explicitly, as the mean separation between the sensing target region and the communication user region increases, maintaining both objectives simultaneously becomes increasingly difficult.
@@ VI Conclusions
A novel deployment framework was proposed for ISAC networks that addresses the fundamental tradeoff between S&C. Our design ensures both localization coverage and communication performance, guaranteeing high-precision localization and high-throughput data transmission across the entire service region. We analyzed the complexity of cooperative sensing among multiple BSs from a localization-performance coverage perspective, ensuring that the CRLB requirements are met uniformly across the region. By identifying key invariance properties of the optimal deployment, such as shift, rotation, and scaling invariance, we developed a low-complexity algorithmic framework for ISAC network planning. We formulated the joint sensing-communication optimization as a structured problem amenable to the MM principle, resulting in an MM-based algorithm that reliably converges to high-quality solutions at a low computational cost per iteration. Simulation results demonstrated that the proposed framework significantly enhances sensing coverage, while maintaining communication throughput with minimal performance degradation. This work provides insights for designing practical large-scale ISAC networks. Exploring joint resource scheduling and BS placement in dynamic scenarios offers valuable avenues for future research. Moreover, extending the area CRLB to exploit informative priors, using Bayesian CRLB formulations, merits further investigation.